OpenStudy (anonymous):

how do solve 48=-3(t-5)*squared* +75

OpenStudy (lgbasallote):

isolate t first... 48 - 75 = -3(t-5)^2 ... -27 = -3(t-5)^2 -27/-3 = (t-5)^2 9 = (t-5)^2 take the square root... 3 = t- 5 t = 8

OpenStudy (anonymous):

thank you !!!!